# Understanding Knowledge Hijack Mechanism in In-context Learning through Associative Memory

### v1

@@ C.3 Induction head without positional encoding
Appendix D Experimental setup
D.1 Oversight of in-context knowledge
@@ D.3 About datasets
Appendix E Other experimental results and analyses
(a) Attention patterns of $\operatorname{TF}_{\text{APE}}$
(b) Attention patterns of $\operatorname{TF}_{\text{RPE}}$
Figure 4: Two-layer transformers with APE and RPE that are trained on sequences of length $128$ show different first-layer attention patterns for length-$256$ sequences. The previous token head of $\operatorname{TF}_{\text{APE}}$ do not function well for positions $t>100$, while $\operatorname{TF}_{\text{RPE}}$ attends to previous tokens regardless of the positions.
E.1 Generalization to longer sequence
acc. ($t<128$) acc. ($t<256$) score ($t<128$) score ($t<256$)
$\operatorname{TF}_{\text{APE}}$ $0.8278\pm 0.0000$ $0.7925\pm 0.0000$ $0.5460\pm 0.0007$ $0.1979\pm 0.0001$
$\operatorname{TF}_{\text{RPE}}$ $0.8727\pm 0.0000$ $0.8974\pm 0.0000$ $0.9513\pm 0.0000$ $0.9337\pm 0.0001$
Table 3: The evaluation results under the change in sequence length from $128$ during training to $256$ during evaluation. Transformer with APE degrades its accuracy and attention score as the sequence length grows, while transformer with RPE shows a stable accuracy and score.
To evaluate the generalization ability of $\operatorname{TF}_{\text{ape}}$ and $\operatorname{TF}_{\text{rpe}}$, we trained these two types of transformers with sequences of length $128$ generated from the bigram model. Details regarding the training procedure are provided in the Appendix D. We measured the accuracy of $\operatorname{TF}_{\text{APE}}$ and $\operatorname{TF}_{\text{RPE}}$ with respect to the second and subsequent occurrences of the output tokens. In addition, We present the memory recall to examine the generalization.
@@ E.1 Generalization to longer sequence
In fact, the attention pattern heatmaps in Fig. 4 in Appendix E tell us that the attention is no longer directed to the previous token after $t>100$ for $\operatorname{TF}_{\text{APE}}$, while $\operatorname{TF}_{\text{RPE}}$ continues to attend to the previous token. These experimental results match our theoretical result, where RPE is better suited for length generalization. It can be seen from Fig. 4(b) that the first-layer attention head of $\operatorname{TF}_{\text{RPE}}$ sometimes attends to the current token. This phenomenon is also explainable from the proof of Theorem 3, which shows that the attention matrix $W_{K}^{1}$ not only learns the associations between $w_{E}(v)$ and $r_{-1}$, but also the pair of $w_{E}(v)$ and $r_{0}$.
E.2 Discussion of collision of context information
We discuss the result shown in Fig. 3. The figure illustrates that it is more likely to output global knowledge when the number of $B_{1}$ and $B_{2}$ are the same, or when the number of $B_{1}$ or $B_{2}$ is almost maximum. Using Proposition 2, we can explain these phenomena by considering inequalities using the global knowledge $\log{\pi_{b}(\cdot\mid A)}$ and the strength of in-context knowledge $\tau_{3}$. Using Proposition 2, we can explain these phenomena as follows. If we have the relationships $\log{\pi_{b}(B_{*}\mid A)}+\tau_{3}>\log{\pi_{b}(B\mid A)}>\log{\pi_{b}(B_{*}\mid A)}+\tau_{3}/2$, where $B_{*}$ represents $B_{1}$ and $B_{2}$, the model outputs global knowledge $B$ when $AB_{1}$ and $AB_{2}$ occur equally frequently in the prompt. On the other hand, if we have $\log{\pi_{b}(B_{2}\mid A)}+\tau_{3}>\log{\pi_{b}(B\mid A)}>\log{\pi_{b}(B_{1}\mid A)}+\tau_{3}$ and $\log{\pi_{b}(B\mid A)}>\log{\pi_{b}(B_{2}\mid A)}$, the prediction of the model changes to $B_{2}$ to $B$ as the number of $B_{1}$ increases.

### v2

@@ D.3 Induction head without positional encoding
Appendix E Experimental setup
E.1 Oversight of in-context knowledge
@@ E.3 About datasets
Appendix F Other experimental results and analyses
F.1 Generalization to longer sequence
To evaluate the generalization ability of $\operatorname{TF}_{\text{ape}}$ and $\operatorname{TF}_{\text{rpe}}$, we trained these two types of transformers with sequences of length $128$ generated from the bigram model. Details regarding the training procedure are provided in the Appendix E. We measured the accuracy of $\operatorname{TF}_{\text{APE}}$ and $\operatorname{TF}_{\text{RPE}}$ with respect to the second and subsequent occurrences of the output tokens. In addition, We present the memory recall to examine the generalization.
@@ F.1 Generalization to longer sequence
In fact, the attention pattern heatmaps in Fig. 3 in Appendix F tell us that the attention is no longer directed to the previous token after $t>100$ for $\operatorname{TF}_{\text{APE}}$, while $\operatorname{TF}_{\text{RPE}}$ continues to attend to the previous token. These experimental results match our theoretical result, where RPE is better suited for length generalization. It can be seen from Fig. 3(b) that the first-layer attention head of $\operatorname{TF}_{\text{RPE}}$ sometimes attends to the current token. This phenomenon is also explainable from the proof of Theorem 3, which shows that the attention matrix $W_{K}^{1}$ not only learns the associations between $w_{E}(v)$ and $r_{-1}$, but also the pair of $w_{E}(v)$ and $r_{0}$.
F.2 Discussion of collision of context information
We discuss the result shown in Fig. 1(b). The figure illustrates that it is more likely to output global knowledge when the number of $B_{1}$ and $B_{2}$ are the same, or when the number of $B_{1}$ or $B_{2}$ is almost maximum. Using Proposition 2, we can explain these phenomena by considering inequalities using the global knowledge $\log{\pi_{b}(\cdot\mid A)}$ and the strength of in-context knowledge $\tau_{3}$. Using Proposition 2, we can explain these phenomena as follows. If we have the relationships $\log{\pi_{b}(B_{*}\mid A)}+\tau_{3}>\log{\pi_{b}(B\mid A)}>\log{\pi_{b}(B_{*}\mid A)}+\tau_{3}/2$, where $B_{*}$ represents $B_{1}$ and $B_{2}$, the model outputs global knowledge $B$ when $AB_{1}$ and $AB_{2}$ occur equally frequently in the prompt. On the other hand, if we have $\log{\pi_{b}(B_{2}\mid A)}+\tau_{3}>\log{\pi_{b}(B\mid A)}>\log{\pi_{b}(B_{1}\mid A)}+\tau_{3}$ and $\log{\pi_{b}(B\mid A)}>\log{\pi_{b}(B_{2}\mid A)}$, the prediction of the model changes to $B_{2}$ to $B$ as the number of $B_{1}$ increases.